# Solving nonconvex Hamilton–Jacobi–Isaacs equations with PINN-based policy iteration

### v1

@@ 1 Introduction
The rest of the paper is organized as follows. In Section 2, we breifly review the formulation of Hamilton–Jacobi–Isaacs equations in the context of differential games. Section 3 presents the proposed PINN-based policy iteration framework, detailing the policy evaluation and improvement steps, the associated training procedure, and a theoretical analysis of the method, including a convergence guarantee under suitable regularity and ellipticity conditions. Section 4 presents the results of extensive numerical experiments on benchmark problems in multiple dimensions, highlighting the accuracy, scalability, and robustness of the proposed approach. Finally, Section 5 concludes the paper and discusses potential directions for future research.
2 Stochastic Differential Games and Hamilton–Jacobi–Issacs equations
@@
$\begin{cases}\mathrm{d}X(s)&=f(X(s),a(s),b(s))\mathrm{d}s+\sigma(s,X(s))\mathrm{d}W_{s},\quad\text{for}\quad s\geq t,\\ X(t)&=x,\end{cases}$ (2.1)
where $W_{s}\in\mathbb{R}^{d}$ denotes a standard $d$-dimensional Brownian motion and the functions $f:\mathbb{R}^{d}\times A\times B\to\mathbb{R}^{d}$ and $\sigma:[0,T]\times\mathbb{R}^{d}\to\mathbb{R}^{d\times d}$ represent the drift and diffusion coefficients, respectively. Let $\mathcal{F}:=(\mathcal{F}_{s})_{s\geq 0}$ be the filtration generated by $(W_{s})_{s\geq 0}$, and for $s\in[t,T]$, set $a\in\mathcal{A}_{t}$ and $b\in\mathcal{B}_{t}$ where
@@
$J(t,x;a,b)=\mathbb{E}[\int_{t}^{T}c(X(s),a(s),b(s))\mathrm{d}s+g(X_{T})],$
@@
Under standard regularity assumptions, such as Lipschitz continuity and boundedness of $f$, $c$, and $\sigma$ in both $x$ and $t$, and uniform ellipticity of the matrix $\sigma(t,x)\sigma^{\top}(t,x)$, the value function $v(t,x)$ satisfies the dynamic programming principle and is characterized as the unique viscosity solution of the Hamilton–Jacobi–Isaacs (HJI) equation
@@ 3.1 Policy iteration for HJI equations
$\|f\|_{p}:=(\int_{\Omega}|f|^{p})^{1/p}\mathrm{d}x,\quad\text{for}\quad p\in(0,\infty],$
where $|f|$ denotes the pointwise Euclidean norm of $f$. We denote by $\mathrm{Lip}(\Omega)$ the set of Lipschitz continuous functions on $\Omega$, and by $C(\Omega)$ the set of continuous functions.
@@
$(a,b)\mapsto\arg\min_{a\in A}\arg\max_{b\in B}L(t,x,\nabla_{x}v)(a,b)$
@@
In contrast, we assume that the coefficients $f$, $c$, and $\sigma$ are Hölder continuous in $(t,x)$ and that the terminal cost $g\in C^{2+\beta}(\mathbb{R}^{d})$, we obtain, via Schauder theory, that each value function $v_{n}$ belongs to $C^{2+\beta}$ in space variable $x$, and hence, has Lipschitz continuous gradients $\nabla_{x}v_{n}$ with respect to $x$.
@@
  • The diffusion coefficient $\sigma\in C([0,T]\times\mathbb{R}^{d};\mathbb{R}^{d\times d})$ is uniformly elliptic: there exists $\lambda>0$ such that
    $\sigma(t,x)\sigma(t,x)^{\top}\succeq\lambda I_{d}\quad\text{for all}\quad(t,x)\in[0,T]\times\mathbb{R}^{d}.$
@@
For $p\in\mathbb{R}^{d}$, set $F_{p}:=(\nabla_{\mathbf{a}}L_{p},-\nabla_{\mathbf{b}}L_{p})$. Since $L_{p}$ is $\mu_{A}$–strongly convex in $\mathbf{a}$ and $\mu_{B}$–strongly concave in $\mathbf{b}$, $F_{p}$ is $\mu$‑strongly monotone with $\mu:=\min\{\mu_{A},\mu_{B}\}$:
@@
$v_{n}(t,x)=\mathbb{E}[g(X_{T}^{t,x})+\int_{t}^{T}c_{n}(s,X_{s}^{t,x})\mathrm{d}s,$
@@
$|v_{n}(t,x)|\leq\mathbb{E}[|g(X_{T}^{t,x})|]+\mathbb{E}[\int_{t}^{T}|c_{n}(s,X_{s}^{t,x})|\mathrm{d}s]\leq\|g\|_{\infty}+T\|c_{n}\|_{\infty}.$
@@
$v_{n}(t,z)=\mathbb{E}[g(X^{t,z}_{T})+\int_{t}^{T}c_{n}(s,X^{t,z}_{s})\mathrm{d}s],\quad z\in\{x,y\}.$
@@
$|v_{n}(t,x)-v_{n}(t,y)|$ $\leq\mathbb{E}[|g(X^{t,x}_{T})-g(X^{t,y}_{T})|]+\mathbb{E}[\int_{t}^{T}|c_{n}(s,X^{t,x}_{s})-c_{n}(s,X^{t,y}_{s})|\mathrm{d}s]$
@@
$v_{n}(t,x)=\mathbb{E}[v_{n}(t+h,X^{t,x}_{t+h})+\int_{t}^{t+h}c_{n}(s,X^{t,x}_{s})\mathrm{d}s].$
@@
where $C$ depends only on $\|c\|_{\infty}$ and $\|\sigma\|_{\infty}$, we obtain
$|v_{n}(t+h,x)-v_{n}(t,x)|\leq C\sqrt{h},$
@@
$\mathcal{L}_{n}:=L(t,x,\nabla_{x}v_{n+1})(\pi_{n+1})-L(t,x,\nabla_{x}v_{n})(\pi_{n}).$
Since
$\mathcal{L}_{n}$ $=\underbrace{L(t,x,\nabla_{x}v_{n+1})(\pi_{n+1})-L(t,x,\nabla_{x}v_{n})(\pi_{n+1})L(t,x,\nabla_{x}v_{n})(\pi_{n+1})-L(t,x,\nabla_{x}v_{n})(\pi_{n})}_{=:\mathrm{II}}.$
Clearly, $|\mathrm{I}|\leq L_{u}|\nabla\delta_{n}|$. To bound $\mathrm{II}$, let us recall the optimality condition of the Hamiltonian, which is
@@
$\begin{split}|\mathrm{II}|&=|H(t,x,p_{n})-H(t,x,p_{n-1})+H(t,x,p_{n-1})-L(t,x,p_{n})(\pi_{n})|\\ &\leq|H(t,x,p_{n})-H(t,x,p_{n-1})|+|L(t,x,p_{n-1})(\pi_{n})-L(t,x,p_{n})(\pi_{n})|\\ &\leq 2L_{u}|p_{n}-p_{n-1}|,\end{split}$
@@
$|\mathcal{L}_{n}|\leq 2L_{u}(|\nabla_{x}\delta_{n}|+|\nabla_{x}\delta_{n-1}|).$ (3.2)
Subtracting the two policy evaluation equations for $v_{n+1}$ and $v_{n}$, multiplying by $\delta_{n}$ and integrating in space gives
@@
$2\int_{\mathbb{R}^{d}}|\mathcal{L}_{n}\delta_{n}|\mathrm{d}x\leq 4L_{u}(\|\nabla_{x}\delta_{n}(t,\cdot)\|_{2}\|\delta_{n}(t,\cdot)\|_{2}+\|\nabla_{x}\delta_{n-1}(t,\cdot)\|_{2}\|\delta_{n}(t,\cdot)\|_{2}).$
Invoking Young’s inequality $ab\leq\tfrac{\eta}{2}a^{2}+\tfrac{1}{2\eta}b^{2}$ with $\eta=\lambda/(4L_{u})$ gives
$2\int_{\mathbb{R}^{d}}|\mathcal{L}_{n}\delta_{n}|\mathrm{d}x\leq\frac{\lambda}{2}\|\nabla_{x}\delta_{n}(t,\cdot)\|_{2}^{2}+\frac{\lambda}{2}\|\nabla_{x}\delta_{n-1}(t,\cdot)\|_{2}^{2}+\frac{16L_{u}^{2}}{\lambda}\|\delta_{n}(t,\cdot)\|_{2}^{2}.$ (3.4)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\|\delta_{n}(t,\cdot)\|_{2}^{2}+\frac{\lambda}{2}\|\nabla_{x}\delta_{n}(t,\cdot)\|_{2}^{2}\leq\frac{16L_{u}^{2}}{\lambda}\|\delta_{n}(t,\cdot)\|_{2}^{2}+\frac{\lambda}{2}\|\nabla_{x}\delta_{n-1}(t,\cdot)\|_{2}^{2}.$ (3.5)
@@
$E_{n}^{2}+\frac{\lambda}{2}\int_{0}^{T}\|\nabla_{x}\delta_{n}(t,\cdot)\|_{2}^{2}\mathrm{d}t\leq\frac{16L_{u}^{2}T}{\lambda}E_{n}^{2}+\frac{\lambda T}{2}F_{n-1}^{2},$
@@
where $\mu=1-\frac{16L_{u}^{2}T}{\lambda}$. Since $\delta_{n}(0,x)=0$, we invoke the gradient bound of $\delta_{n}$ demonstrated in [6], we have
@@ 3.2 PINN-Based Policy Iteration
1:collocation sets $\{(t_{j},x_{j})\}$, terminal set $\{x_{k}^{T}\}$, tolerance $\mathrm{tol}$, number of policy updates $M$, number of training epochs per policy update $E$
@@
Suppose Assumption 1 hold, and let $\{\tilde{v}_{n},\tilde{\alpha}_{n},\tilde{\beta}_{n}\}_{n\geq 0}$ be generated by Algorithm 2. If $\|\mathcal{R}_{n}\|_{2}=p_{n}$ for
$\mathcal{R}_{n}:=\partial_{t}\tilde{v}_{n}+L(t,x,\tilde{\alpha}_{n}\tilde{\beta}_{n})+\tfrac{1}{2}\mathrm{Tr}(\sigma\sigma^{\top}D^{2}\tilde{v}_{n}),$
@@
where $(\alpha_{n},\beta_{n})$ the exact feedback pair obtained from $v_{n-1}$.
@@
$\begin{cases}\partial_{t}B_{n}+\tfrac{1}{2}\mathrm{Tr}(\sigma\sigma^{\top}D^{2}B_{n})=:-\mathcal{L}_{n},\quad\text{in}\quad(0,T]\times\mathbb{R}^{d},\\ B_{n}(0,\cdot)=0,\quad\text{on}\quad\mathbb{R}^{d}.\end{cases}$
@@
$\begin{split}\sup_{t\in[0,T]}\|B_{n}(t,\cdot)\|_{2}&\leq C\|\mathcal{L}_{n}\|_{2}\\ &\leq C\underbrace{\sup_{t\in[0,T]}\|(\nabla_{x}(\tilde{v}_{n-1}(t,\cdot)-v_{n-1}(t,\cdot))\|_{2}}_{=:F_{n-1}})\end{split}$
@@
This result highlights a key advantage of the proposed framework: unlike black-box direct PINN solvers, the iterative structure permits explicit $L^{2}$ error estimates with provable rates. This is particularly important in nonconvex settings, where bounding the solution error is otherwise analytically intractable.

### v2

@@ 1 Introduction
The rest of the paper is organized as follows. In Section 2, we briefly review the formulation of Hamilton–Jacobi–Isaacs equations in the context of differential games. Section 3 presents the proposed PINN-based policy iteration framework, detailing the policy evaluation and improvement steps, the associated training procedure, and a theoretical analysis of the method, including a convergence guarantee under suitable regularity and ellipticity conditions. Section 4 presents the results of extensive numerical experiments on benchmark problems in multiple dimensions, highlighting the accuracy, scalability, and robustness of the proposed approach. Finally, Section 5 concludes the paper and discusses potential directions for future research.
2 Stochastic Differential Games and Hamilton–Jacobi–Isaacs equations
@@
$\begin{cases}\mathrm{d}X(s)&=f(s,X(s),a(s),b(s))\mathrm{d}s+\sigma(s,X(s))\mathrm{d}W_{s},\quad\text{for}\quad s\geq t,\\ X(t)&=x,\end{cases}$ (2.1)
where $W_{s}\in\mathbb{R}^{d}$ denotes a standard $d$-dimensional Brownian motion and the functions $f:[0,T]\times\mathbb{R}^{d}\times A\times B\to\mathbb{R}^{d}$ and $\sigma:[0,T]\times\mathbb{R}^{d}\to\mathbb{R}^{d\times d}$ represent the drift and diffusion coefficients, respectively. Let $\mathcal{F}:=(\mathcal{F}_{s})_{s\geq 0}$ be the filtration generated by $(W_{s})_{s\geq 0}$, and for $s\in[t,T]$, set $a\in\mathcal{A}_{t}$ and $b\in\mathcal{B}_{t}$ where
@@
$J(t,x;a,b)=\mathbb{E}\left[\int_{t}^{T}c(s,X(s),a(s),b(s))\mathrm{d}s+g(X_{T})\right],$
@@
Under standard regularity assumptions, such as Lipschitz continuity and boundedness of $f$, $c$, and $\sigma$ in both $t$ and $x$, and uniform ellipticity of the matrix $\sigma\sigma^{\top}(t,x)$, the value function $v(t,x)$ satisfies the dynamic programming principle and is characterized as the unique viscosity solution of the Hamilton–Jacobi–Isaacs (HJI) equation
@@ 3.1 Policy iteration for HJI equations
$\|f\|_{p}:=\left(\int_{\Omega}|f|^{p}\mathrm{d}x\right)^{1/p},\quad\text{for}\quad p\in(0,\infty],$
where $|f|$ denotes the pointwise Euclidean norm of $f$. We say $f\in C^{k,\beta}(\Omega)$ for $k\in\mathbb{N}$ and $\beta\in(0,1)$ if all partial derivatives $D^{\alpha}f$ of order $|\alpha|\leq k$ exist and are continuous on $\Omega$, and for all multi-indices $\alpha$ with $|\alpha|=k$,
$[D^{\alpha}f]_{C^{0,\beta}(\Omega)}:=\sup_{x\neq y\in\Omega}\frac{|D^{\alpha}f(x)-D^{\alpha}f(y)|}{|x-y|^{\beta}}<\infty.$
We write $C^{\beta}(\Omega):=C^{0,\beta}(\Omega)$ for the space of $\beta$-Hölder continuous functions, and $C(\Omega)$ for the space of continuous functions.
@@
$(\textbf{a},\textbf{b})\mapsto\arg\min_{\textbf{a}\in A}\arg\max_{\textbf{b}\in B}L(t,x,\nabla_{x}v)(\textbf{a},\textbf{b})$
@@
In contrast, we assume that the coefficients $f$, $c$, and $\sigma$ are Hölder continuous in $(t,x)$ and that the terminal cost $g\in C^{2+\beta}(\mathbb{R}^{d})$, we obtain, via Schauder theory, that each value function $v_{n}$ belongs to $C^{2+\beta}$ in sptial variable $x$, and hence, has Lipschitz continuous gradients $\nabla_{x}v_{n}$ with respect to $x$.
@@
  • The diffusion coefficient $\sigma\sigma^{\top}$ is uniformly elliptic: there exists $\lambda>0$ such that
    $\sigma\sigma(t,x)^{\top}\succeq\lambda I_{d}\quad\text{for all}\quad(t,x)\in[0,T]\times\mathbb{R}^{d}.$
@@
For $p\in\mathbb{R}^{d}$, set $F_{p}:=(\nabla_{\mathbf{a}}L_{p},-\nabla_{\mathbf{b}}L_{p})$ for $L_{p}:=L(t,x,p)(\textbf{a},\textbf{b})$. Since $L_{p}$ is $\mu_{A}$–strongly convex in $\mathbf{a}$ and $\mu_{B}$–strongly concave in $\mathbf{b}$, $F_{p}$ is $\mu$‑strongly monotone with $\mu:=\min\{\mu_{A},\mu_{B}\}$:
@@
$v_{n}(t,x)=\mathbb{E}\left[g(X_{T}^{t,x})+\int_{t}^{T}c_{n}(s,X_{s}^{t,x})\mathrm{d}s\right],$
@@
$|v_{n}(t,x)|\leq\mathbb{E}[|g(X_{T}^{t,x})|]+\mathbb{E}\left[\int_{t}^{T}|c_{n}(s,X_{s}^{t,x})|\mathrm{d}s\right]\leq\|g\|_{\infty}+T\|c_{n}\|_{\infty}.$
@@
$v_{n}(t,z)=\mathbb{E}\left[g(X^{t,z}_{T})+\int_{t}^{T}c_{n}(s,X^{t,z}_{s})\mathrm{d}s\right],\quad z\in\{x,y\}.$
@@
$|v_{n}(t,x)-v_{n}(t,y)|$ $\leq\mathbb{E}[|g(X^{t,x}_{T})-g(X^{t,y}_{T})|]+\mathbb{E}\left[\int_{t}^{T}|c_{n}(s,X^{t,x}_{s})-c_{n}(s,X^{t,y}_{s})|\mathrm{d}s\right]$
@@
$v_{n}(t,x)=\mathbb{E}\left[v_{n}(t+h,X^{t,x}_{t+h})+\int_{t}^{t+h}c_{n}(s,X^{t,x}_{s})\mathrm{d}s\right].$
@@
$\begin{split}\mathcal{L}_{n}:&=L(t,x,\nabla_{x}v_{n+1})(\pi_{n+1})-L(t,x,\nabla_{x}v_{n})(\pi_{n}).\\ &=\underbrace{L(t,x,\nabla_{x}v_{n+1})(\pi_{n+1})-L(t,x,\nabla_{x}v_{n})(\pi_{n+1})}_{=:\mathrm{I}}+\underbrace{L(t,x,\nabla_{x}v_{n})(\pi_{n+1})-L(t,x,\nabla_{x}v_{n})(\pi_{n})}_{=:\mathrm{II}}.\end{split}$
Clearly, $|\mathrm{I}|\leq\|f\|_{\infty}|\nabla\delta_{n}|$. To bound $\mathrm{II}$, let us recall the optimality condition of the Hamiltonian, which is
@@
$\begin{split}|\mathrm{II}|&=|H(t,x,p_{n})-H(t,x,p_{n-1})+H(t,x,p_{n-1})-L(t,x,p_{n})(\pi_{n})|\\ &\leq|H(t,x,p_{n})-H(t,x,p_{n-1})|+|L(t,x,p_{n-1})(\pi_{n})-L(t,x,p_{n})(\pi_{n})|\\ &\leq 2\|f\|_{\infty}|p_{n}-p_{n-1}|,\end{split}$
@@
$|\mathcal{L}_{n}|\leq 2\|f\|_{\infty}(|\nabla_{x}\delta_{n}|+|\nabla_{x}\delta_{n-1}|).$ (3.2)
Subtracting the two policy evaluation equations for $v_{n+1}$ and $v_{n}$, multiplying by $\delta_{n}$ and integrating over $\mathbb{R}^{d}$ gives
@@
$2\int_{\mathbb{R}^{d}}|\mathcal{L}_{n}\delta_{n}|\mathrm{d}x\leq 4\|f\|_{\infty}(\|\nabla_{x}\delta_{n}(t,\cdot)\|_{2}\|\delta_{n}(t,\cdot)\|_{2}+\|\nabla_{x}\delta_{n-1}(t,\cdot)\|_{2}\|\delta_{n}(t,\cdot)\|_{2}).$
Invoking Young’s inequality $ab\leq\tfrac{\eta}{2}a^{2}+\tfrac{1}{2\eta}b^{2}$ with $\eta=\lambda/(4\|f\|_{\infty})$ gives
$2\int_{\mathbb{R}^{d}}|\mathcal{L}_{n}\delta_{n}|\mathrm{d}x\leq\frac{\lambda}{2}\|\nabla_{x}\delta_{n}(t,\cdot)\|_{2}^{2}+\frac{\lambda}{2}\|\nabla_{x}\delta_{n-1}(t,\cdot)\|_{2}^{2}+\frac{16\|f\|_{\infty}^{2}}{\lambda}\|\delta_{n}(t,\cdot)\|_{2}^{2}.$ (3.4)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\|\delta_{n}(t,\cdot)\|_{2}^{2}+\frac{\lambda}{2}\|\nabla_{x}\delta_{n}(t,\cdot)\|_{2}^{2}\leq\frac{16\|f\|_{\infty}^{2}}{\lambda}\|\delta_{n}(t,\cdot)\|_{2}^{2}+\frac{\lambda}{2}\|\nabla_{x}\delta_{n-1}(t,\cdot)\|_{2}^{2}.$ (3.5)
@@
$E_{n}^{2}+\frac{\lambda}{2}\int_{0}^{T}\|\nabla_{x}\delta_{n}(t,\cdot)\|_{2}^{2}\mathrm{d}t\leq\frac{16\|f\|_{\infty}^{2}T}{\lambda}E_{n}^{2}+\frac{\lambda T}{2}F_{n-1}^{2},$
@@
where $\mu=1-\frac{16\|f\|_{\infty}^{2}T}{\lambda}$. Since $\delta_{n}(0,x)=0$, we invoke the gradient bound of $\delta_{n}$ demonstrated in [6], we have
@@ 3.2 PINN-Based Policy Iteration
1:collocation sets $\{(t_{j},x_{j})\}$, terminal set $\{x_{k}^{T}\}$, tolerance ($\mathrm{tol})$, number of policy updates $M$, number of training epochs per policy update $E$
@@
Suppose Assumption 1 holds, and let $\{\tilde{v}_{n},\tilde{\alpha}_{n},\tilde{\beta}_{n}\}_{n\geq 0}$ be generated by Algorithm 2. If $\|\mathcal{R}_{n}\|_{2}=p_{n}$ for
$\mathcal{R}_{n}:=\partial_{t}\tilde{v}_{n}+L(t,x,\tilde{\alpha}_{n},\tilde{\beta}_{n})+\tfrac{1}{2}\mathrm{Tr}(\sigma\sigma^{\top}D_{xx}^{2}\tilde{v}_{n}),$
@@
where $(\alpha_{n},\beta_{n})$ is the exact feedback pair obtained from $v_{n-1}$.
@@
$\begin{cases}\partial_{t}B_{n}+\tfrac{1}{2}\mathrm{Tr}(\sigma\sigma^{\top}D_{xx}^{2}B_{n})=:-\mathcal{L}_{n},\quad\text{in}\quad(0,T]\times\mathbb{R}^{d},\\ B_{n}(0,\cdot)=0,\quad\text{on}\quad\mathbb{R}^{d}.\end{cases}$
@@
$\begin{split}\sup_{t\in[0,T]}\|B_{n}(t,\cdot)\|_{2}&\leq C\|\mathcal{L}_{n}\|_{2}\\ &\leq C\underbrace{\sup_{t\in[0,T]}\|\nabla_{x}(\tilde{v}_{n-1}(t,\cdot)-v_{n-1}(t,\cdot))\|_{2}}_{=:F_{n-1}})\end{split}$
@@
This result highlights a key advantage of the proposed framework: unlike black-box direct PINN solvers, the iterative structure permits explicit $L^{2}$-error estimates with provable rates. This is particularly important in nonconvex settings, where bounding the solution error is otherwise analytically intractable.